OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = 1/3x – 5/3; (–4, 3)

OpenStudy (anonymous):

y = –3x + 15 y = –3x – 9 y = –4x + 3 y = –1/3x – 11

OpenStudy (anonymous):

i think it is b

OpenStudy (zehanz):

Use the rule: if lines are perpendicular then m1*m2=-1, where m1 and m2 are the slopes of the lines. The slope of your line is 1/3. With the rule you can work out the other slope.

OpenStudy (zehanz):

Once you've got that other slope, you have this equation for tha line: y=m2*x + b. To get b, use the given point: (-4,3), so set x=-4 and y=3.

OpenStudy (zehanz):

If you do this calculation, you will see b is wrong...

OpenStudy (anonymous):

im honestly so lost on this one.

OpenStudy (zehanz):

1/3 * -3 = -1, so m2 = -3. Equation: y=-3x+b. Point on line: (-4,3). Set x=-4, y=3: 3=-3*-4+b so 3 = 12 + b b = 3 -12 = -9. Equation y = -3 x -9 Answer b IS right , I was WRONG! (Sorry...)